1. A 17.0 kg curling stone is initially moving at 4.5 m/s [forward] along the ice. The coefficient of kinetic friction for the stone and the ice is 0.014. What is the final velocity when it reaches its target that is 27.5 m away
  2. Provide a real-life example to briefly explain Newton's Third Law.
  3. A 17.0 kg curling stone is initially moving at 4.5 m/s [forward] along the ice. The coefficient of kinetic friction for the stone and the ice is 0.014. How many arrows will the FBD of the curling stone have? Explain what each arrow represent, and which direction it will point to.
  4. Harry is hanging on to a rope on the side of a cliff. His friends on the top of the cliff is holding on to the other end of the rope without pulling him up. For Harry's FBD, how many arrows are there, and what does each arrow represent?
  5. Continuing from the previous question about Harry hanging on to a rope. For Harry's friends' FBD, how many arrows are there, and what does each arrow represent?
1
Expert's answer
2021-07-16T08:51:54-0400

1)


"v^2-v_i^2=-2\mu g s\\v^2-4.5^2=-2(0.014)(9.8)(27.5)\\v=3.6\frac{m}{s}"

2) For example, when you jump, your legs apply a force to the ground, and the ground applies and equal and opposite reaction force that propels you into the air. 

3) 3 arrows: normal force (upwards), friction force (backwards) and weight (downwards).

4) 2 arrows: pushing force and weight.

5) 3 arrows: normal force, tension and weight.